Physical Chemistry

18g of glucose is added to 178.2g of water. The vapour pressure of this aqueous at 100^0C is:
a.76torr b.752.40torr

Tanusri Bose
7 Years agoGrade

1 Answer

Arun
7 Years ago

Molecular mass of water = 2×1 + 1×16 = 18 g

For 178.2 g water, nA = 9.9

Molecular mass of glucose = 6×12 + 12×1 + 6×16 = 180 g

For 18 g glucose, nB = 0.1

XB = 0.1/(0.1+9.9) = 0.01

XA = 0.99

For lowering of vapour pressure,

P = p0AXA = p0A(1 – XB)

P = 760(1 – 0.01)

= 760 - 7.6

= 752.4 torr